Supto:

An airplane taking off from a landing field has a run of 500m. If it starts from rest, moves with constant acceleration, and makes the runin 30s, with what velocity in m.s-ยน did it tak e off ?

supie:

Is there choices?

supie:

Do you know if theres a formula for this?

supie:

If you find the formula that needs to be used, then you could plug in the numbers and solve the equation.

SemiDefinite:

SUVAT u = 0 s = 500 t = 30 a = constant, ergo SUVAT v <-- what you want to know \(s = ut + \dfrac{1}{2} a t^2 \implies 500 = \dfrac{1}{2} a (30)^2\) \(\therefore a = \dfrac{2 \times 500}{30^3}= \dfrac{10}{9} m s^{-2}\) \(v = u + at = \dfrac{10}{9} \times 30 = 30~m/s\) #Utopia